OpenStudy (anonymous):

In Madurodam, one of the models of a windmill is 1.2 meters tall and 0.8 meters wide. The width of the actual windmill is 20 meters. What is its height? A. 34 meters B. 30 meters C. 20.4 meters D. 24 meters The ratio of the width of the model windmill to the width of the real windmill provides the scale on which the city of Madurodam is built. The scale of the city is 1: ___. (Enter only the number.)

OpenStudy (anonymous):

B 30 meters

OpenStudy (anonymous):

1:25

OpenStudy (anonymous):

\[1.2h = 0.8w \rightarrow h = (3/2)w \rightarrow h = (3/2)(20) \rightarrow h = 30\] scale = model height : actual height scale = 1.2 : 30 scale = 1 : 25